# Rare and exclusive few-body decays of the Higgs, Z, W bosons, and the top quark

### v3

@@ III.2 Exclusive Higgs decays into a gauge boson plus a meson
Since it is extremely difficult to experimentally access the Yukawa couplings to first-generation ($\rm q=u,d,s$) and second-generation (c) quarks, due to the smallness of the $\rm H\to\mathrm{q}\overline{\mathrm{q}},\mathrm{c}\overline{\mathrm{c}}$ decay widths, and the very large QCD jet backgrounds at the LHC, it has been proposed to constrain those couplings via rare exclusive decays into a photon (or a gauge boson) plus a vector meson [11, 12, 13, 14, 15]. The relevant Feynman diagrams for the $\rm H\to V+M$ (where $\rm V=Z,W^{\pm},\gamma$, and M = meson) decays in the SM are shown in Fig. 4. The first diagram represents LO amplitudes where the Higgs boson directly couples to a quark-antiquark pair that radiates a gauge boson and forms the mesonic bound state. The second diagram depicts indirect contributions to the decay amplitude, where the scalar boson decays first into two gauge bosons, one of which transforms (offshell) into a hadron state via $\rm V^{*}\to\mathrm{q}\overline{\mathrm{q}}$. The third diagram shows the radiative FCNC decay into a photon plus a flavoured meson, through a double W loop. The indirect diagram provides the dominant contribution to the decay rates due to the smallness of the Yukawa couplings to the first- and second-generation quarks, and the largest sensitivity to the Higgs-quark coupling comes from the (destructive) interference of the two amplitudes. Thus, for example, the $\rm H\to\gamma+\mathrm{J}/\psi,\gamma+\phi$ modes allow direct access to the flavour-diagonal coupling of the Higgs boson to the charm and strange quarks, respectively, while the $\rm H\to\gamma+\rho,\gamma+\omega$ decays can probe the Higgs couplings to up and down quarks. The radiative decays $\rm H\to\gamma+M$, where $\rm M=K^{*0},D^{*0},B_{s}^{*0},B_{d}^{*0}$ – which in the SM can only proceed through a virtual W boson because the photon (and Z) splitting preserves flavour, or through the direct process in BSM scenarios with flavour-changing Higgs decays – provide possibilities to probe BSM flavour-violating q-q’ Higgs couplings [13, 91].
@@
where $Q_{\mathrm{q}}$ is the electric charge of the constituent quarks, and $m_{\mathrm{M}}$ the mass of the meson. Corrections due to the offshellness of the photon and to the contribution of the $\rm H\to\gamma\,Z^{*}$ process are suppressed by $m^{2}_{\mathrm{M}}/m^{2}_{\mathrm{H}}$ and $m^{2}_{\mathrm{M}}/(m^{2}_{\mathrm{Z}}-m^{2}_{\mathrm{M}})$, respectively, and hence very small [14]. The direct amplitudes, which are much smaller than the indirect ones except for the heaviest $\rm H\to\gamma+\Upsilon$ decay, have been calculated theoretically by different groups. The hierarchy $m_{\mathrm{M}}\ll m_{\mathrm{H}}$ implies that the vector meson is emitted at very high energy $E_{\mathrm{M}}\gg m_{\mathrm{M}}$ in the Higgs boson rest frame. The constituent partons of the vector meson can thus be described by energetic particles moving collinear to the direction of M, and QCD factorization, either in the SCET or NRQCD incarnation, can be employed to compute them.
@@ III.2.1 Higgs decays into a photon plus a vector meson
Table 4 lists the corresponding theoretical predictions and experimental limits for Higgs decays into a photon plus a vector meson, and Fig. 5 displays them in graphical form. Theoretical $\mathcal{B}$ values are in the range of $\mathcal{O}(10^{-5}$$10^{-9})$, with larger rates for decays into light-quark and charm vector mesons compared to bottomonium due to stronger cancellation of direct-indirect contributions for the latter. Experimental upper bounds have been set for all decays at the LHC [92, 93, 94, 95, 96] except for the $\rm H\to\gamma+D,\gamma+B$ ones, but it seems that evidence for a few of those SM decays will only possible at FCC-ee. The heaviest bottomonium radiative decays will require the number of Higgs bosons produced in a machine like FCC-hh.
Table 4: Compilation of exclusive Higgs decays to a photon plus a meson. For each decay, we provide the predicted branching fraction(s) and the theoretical approach used to compute it, as well as the current experimental upper limit and that conservatively estimated for HL-LHC. The last two columns indicate whether the decay can be produced at FCC-ee/FCC-hh.
@@
$\rm H\to\,\gamma$ $+$ $\rm M$ Branching fraction Framework 2024 HL-LHC FCC-ee FCC-hh
$\rm H\,\to\,\gamma$ + $\rm\rho^{0}$  $(1.68\pm 0.08)$$\times~{}10^{-5}$ SCET+LCDA [14] $<$ $3.7$×$\times~{}10^{-4}$ [97] $\lesssim 5.7\times~{}10^{-5}$
$\rm\omega$  $(1.48\pm 0.08)$$\times~{}10^{-6}$ SCET+LCDA [14] $<$ $5.5$×$\times~{}10^{-4}$ [95] $\lesssim 8.2\times~{}10^{-5}$
$\rm\phi$  $(2.31\pm 0.11)$$\times~{}10^{-6}$ SCET+LCDA [14] $<$ $3.0$×$\times~{}10^{-4}$ [97] $\lesssim 4.5\times~{}10^{-5}$
$\rm J/\psi$  $(3.01\pm 0.15)$$\times~{}10^{-6}$ NRQCD (NLL)+LDME [98] $<$ $2.0$×$\times~{}10^{-4}$ [99] $\lesssim 3.9\times~{}10^{-5}$ [60]
 $\left(2.99^{+0.16}_{-0.15}\right)$$\times~{}10^{-6}$ NRQCD+LCDA [100]
 $(2.95\pm 0.17)$$\times~{}10^{-6}$ SCET+LCDA [14]
$\rm\psi(2S)$  $(1.03\pm 0.06)$$\times~{}10^{-6}$ SCET+LCDA333private communication with author of Ref. [100] [96] $<$ $9.9$×$\times~{}10^{-4}$ [96] $\lesssim 1.4\times~{}10^{-4}$
$\rm\Upsilon(1S)$  $3.0$$\times~{}10^{-8}$ NRQCD (NLO)+LDME [101] $<$ $2.5$×$\times~{}10^{-4}$ [99] $\lesssim 3.8\times~{}10^{-5}$
 $\left(9.97^{+4.04}_{-3.03}\right)$$\times~{}10^{-9}$ NRQCD (NLL)+LDME [98]
 $\left(5.22^{+2.02}_{-1.70}\right)$$\times~{}10^{-9}$ NRQCD+LCDA [100]
 $\left(4.61^{+1.76}_{-1.23}\right)$$\times~{}10^{-9}$ SCET+LCDA [14]
$\rm\Upsilon(2S)$  $1.4$$\times~{}10^{-8}$ NRQCD (NLO)+LDME [101] $<$ $4.2$×$\times~{}10^{-4}$ [99] $\lesssim 6.4\times~{}10^{-5}$
 $\left(2.62^{+1.39}_{-0.91}\right)$$\times~{}10^{-9}$ NRQCD (NLL)+LDME [98]
 $\left(1.42^{+0.72}_{-0.57}\right)$$\times~{}10^{-9}$ NRQCD+LCDA [100]
 $\left(2.34^{+0.76}_{-1.00}\right)$$\times~{}10^{-9}$ SCET+LCDA [14]
$\rm\Upsilon(3S)$  $1.1$$\times~{}10^{-8}$ NRQCD (NLO)+LDME [101] $<$ $3.4$×$\times~{}10^{-4}$ [99] $\lesssim 5.2\times~{}10^{-5}$
 $\left(1.87^{+1.05}_{-0.69}\right)$$\times~{}10^{-9}$ NRQCD (NLL)+LDME [98]
 $\left(9.1^{+4.8}_{-3.8}\right)$$\times~{}10^{-10}$ NRQCD+LCDA [100]
 $\left(2.13^{+0.76}_{-1.12}\right)$$\times~{}10^{-9}$ SCET+LCDA [14]
@@ III.2.2 Higgs decays into a Z boson plus a meson
Higgs decays of the form $\rm H\to Z+M$ involve neutral currents and are very similar to the radiative decays just discussed. However, an important difference with respect to the $\gamma\,+\,$meson decays, is that now since the massive final-state gauge boson can be in a longitudinal polarization state, both pseudoscalar and vector mesons can be produced. Also, the interference terms between the direct and indirect amplitudes are smaller. Although these decays are less useful for measuring the Higgs couplings to light quarks, they probe the important effective H-$\gamma$-Z coupling, and provide independent constraints on the meson distribution amplitudes. In addition, they constitute a background for exotic BSM decays of the Higgs boson into, e.g., a Z boson plus an ALP ($\mathrm{H}\to\mathrm{Z}+a$[102].
@@
Table 5 compiles the theoretical predictions and experimental limits for Higgs decay branching ratios into a Z boson plus a meson, and Fig. 6 presents them in graphical form. All decays have rates in the $10^{-5}$$10^{-6}$ range with small differences across models for the same channel, except for the $\rm H\to Z+\rho$ case where the calculations of [12] and [103] differ by a factor of six, seemingly because the former had neglected the indirect contributions. Experimental upper bounds have been set for a few channels listed in Table 5, and they are in the $\mathcal{O}(10^{-2}$$10^{-3})$ range [^4] We note that one could a priori also derive limits for $\mathcal{B}(\rm H\to Z+\eta_{\mathrm{c}})$ and $\mathcal{B}(\rm H\to Z+\eta_{\mathrm{b}})$ from a recent 95% CL upper limit on the cross section for the $\rm H\to Z(\ell\ell)a(\gamma\gamma)$ process over $m_{a}\approx 1$–30 GeV performed by the CMS Collaboration [102]. However, the upper bound cross sections measured at the $m_{a}=m_{\eta_{\mathrm{c}},\eta_{\mathrm{b}}}\approx 2.9,~{}9.4$ GeV mass points: $\sigma\lesssim 7,$ 2.1 fb, corresponding to $\mathcal{B(\rm H\to Z+a)}\lesssim 0.002,~{}0.62\cdot 10^{-3}$ after normalizing by the Higgs production cross section and correcting for the dilepton decays of the Z boson) are too weak to place any sensible constrain on these $\mathcal{B}$ (namely, they would exclude the nonphysical region $\mathcal{B}(\rm H\to Z+\eta_{\mathrm{c,b}})\gtrsim 10$, dividing them by the corresponding $\mathcal{B}(\eta_{\mathrm{c,b}}\to\gamma\gamma)$ values).. In general, the experimental $\rm H\to Z+M$ limits are about a factor of ten worse than their $\rm H\to\gamma+M$ counterparts because of the extra events lost from the requirement of Z boson identification via dilepton decays (with $\mathcal{B}\approx 3\%$ for each charged lepton pair). Conservatively, the HL-LHC will be able to set upper bounds about 100 times above their expected SM branching fractions, and it appears that experimental evidence will only be possible at FCC-ee for all of them. Bottomonia-plus-Z and $\eta_{\mathrm{c}}$-plus-Z decays have the largest rates, but no upper-limit has been set to date for them.
@@
Table 5: Compilation of exclusive Higgs decays to a Z boson plus a meson. For each decay, we provide the predicted branching fraction(s) and the theoretical approach used to compute it, as well as the current experimental upper limit and that estimated for HL-LHC. The last two columns indicate whether the decay can be produced at FCC-ee/FCC-hh.
@@
$\rm H\to\,Z$ $+$ $\rm M$ Branching fraction Framework 2024 HL-LHC FCC-ee FCC-hh
$\rm H\,\to\,Z$ + $\rm\pi^{0}$  $(2.3\pm 0.1)$$\times~{}10^{-6}$ EFT+NRQM [10]
 $(2.3\pm 0.1)$$\times~{}10^{-6}$ EFT+LCDA [103]
$\rm\eta$  $(8.3\pm 0.9)$$\times~{}10^{-7}$ EFT+LCDA [103]
$\rm\rho^{0}$  $(1.4\pm 0.1)$$\times~{}10^{-5}$ EFT+NRQM [10] $<$ $1.2$×$\times~{}10^{-2}$ [104] $\lesssim 1.8\times~{}10^{-3}$
 $(7.19\pm 0.29)$$\times~{}10^{-6}$ EFT+LCDA [103]
$\rm\omega$  $(1.6\pm 0.1)$$\times~{}10^{-6}$ EFT+NRQM [10]
 $(5.6\pm 0.2)$$\times~{}10^{-7}$ EFT+LCDA [103]
$\rm\eta^{\prime}$  $(1.24\pm 0.13)$$\times~{}10^{-6}$ EFT+LCDA [103]
$\rm\phi$  $(4.2\pm 0.3)$$\times~{}10^{-6}$ EFT+NRQM [10] $<$ $3.6$×$\times~{}10^{-3}$ [104] $\lesssim 5.4\times~{}10^{-4}$
 $(2.42\pm 0.10)$$\times~{}10^{-6}$ EFT+LCDA [103]
$\rm\eta_{c}$  $(1.00\pm 0.01)$$\times~{}10^{-5}$ EFT+LCDA [105]
 $(1.0\pm 0.1)$$\times~{}10^{-5}$ EFT+NRQM [10]
$\rm J/\psi$  $3.4$$\times~{}10^{-6}$ NRQCD (NLO)+LMDE [106] $<$ $1.9$×$\times~{}10^{-3}$ [107] $\lesssim 2.1\times~{}10^{-4}$ [61]
 $(2.3\pm 0.1)$$\times~{}10^{-6}$ EFT+LCDA [103]
 $3.2$$\times~{}10^{-6}$ EFT+NRQM [108]
$\rm\psi(2S)$  $1.5$$\times~{}10^{-6}$ EFT+NRQM [108] $<$ $6.6$×$\times~{}10^{-3}$ [107] $\lesssim 1.0\times~{}10^{-3}$
$\rm\eta_{b}$  $(2.69\pm 0.05)$$\times~{}10^{-5}$ EFT+LCDA [105]
 $\left(4.739^{+0.276}_{-0.244}\right)$$\times~{}10^{-5}$ EFT (NLO)+LCDA [109]
$\rm\Upsilon(1S)$  $1.7$$\times~{}10^{-5}$ NRQCD (NLO)+LMDE [106]
 $(1.54\pm 0.06)$$\times~{}10^{-5}$ EFT+LCDA [103]
 $1.7$$\times~{}10^{-5}$ EFT+NRQM [108]
$\rm\Upsilon(2S)$  $(7.5\pm 0.3)$$\times~{}10^{-6}$ EFT+LCDA [103]
 $8.9$$\times~{}10^{-6}$ EFT+NRQM [108]
$\rm\Upsilon(3S)$  $(5.63\pm 0.24)$$\times~{}10^{-6}$ EFT+LCDA [103]
 $6.7$$\times~{}10^{-6}$ EFT+NRQM [108]

### v4

@@ III.2 Exclusive Higgs decays into a gauge boson plus a meson
Since it is extremely difficult to experimentally access the Yukawa couplings to first-generation ($\rm q=u,d,s$) and second-generation ($\rm q=s,c$) quarks, due to the smallness of the $\rm H\to\mathrm{q}\overline{\mathrm{q}},\mathrm{c}\overline{\mathrm{c}}$ decay widths, and the very large QCD dijet backgrounds at the LHC, it has been proposed to constrain those couplings via rare exclusive decays into a photon (or a gauge boson) plus a vector meson [Bodwin:2013gca, Isidori:2013cla, Kagan:2014ila, Konig:2015qat, Perez:2015lra]. The relevant Feynman diagrams for the $\rm H\to V+M$ (where $\rm V=Z,W^{\pm},\gamma$, and M = meson) decays in the SM are shown in Fig. 4. The first diagram represents LO amplitudes where the Higgs boson directly couples to a quark-antiquark pair that radiates a gauge boson and forms the mesonic bound state. The second diagram depicts indirect contributions to the decay amplitude, where the scalar boson decays first into two gauge bosons, one of which transforms (offshell) into a hadron state via $\rm V^{*}\to\mathrm{q}\overline{\mathrm{q}}$. The third diagram shows the radiative FCNC decay into a gauge boson plus a flavoured meson, through a double W loop. The indirect diagram provides the dominant contribution to the decay rates due to the smallness of the Yukawa couplings to the first- and second-generation quarks, and the largest sensitivity to the Higgs-quark coupling comes from the (destructive) interference between the two amplitudes. Thus, for example, the $\rm H\to\gamma+\mathrm{J}/\psi,\gamma+\phi$ modes allow direct access to the flavour-diagonal coupling of the Higgs boson to the charm and strange quarks, respectively, while the $\rm H\to\gamma+\rho,\gamma+\omega$ decays can probe the Higgs couplings to up and down quarks. The radiative decays $\rm H\to\gamma+M$, where $\rm M=K^{*0},D^{*0},B^{*0},B_{s}^{*0}$ – which in the SM can only proceed through a virtual W boson because the photon (and Z) splitting preserves flavour, or through the direct process in BSM scenarios with flavour-changing Higgs decays – provide possibilities to probe BSM flavour-violating q-q’ Higgs couplings [Kagan:2014ila, Kamenik:2023hvi].
@@
where $Q_{\mathrm{q}}$ is the electric charge of the constituent quarks, and $m_{\mathrm{M}}$ the mass of the meson. Corrections due to the offshellness of the photon and to the contribution of the $\rm H\to\gamma\,Z^{*}$ process are suppressed by $m^{2}_{\mathrm{M}}/m^{2}_{\mathrm{H}}$ and $m^{2}_{\mathrm{M}}/(m^{2}_{\mathrm{Z}}-m^{2}_{\mathrm{M}})$, respectively, and hence very small [Konig:2015qat]. The direct amplitudes, which are much smaller than the indirect ones except for the heaviest $\rm H\to\gamma+\Upsilon$ decay, have been calculated theoretically by different groups. The hierarchy $m_{\mathrm{M}}\ll m_{\mathrm{H}}$ implies that the vector meson is emitted at very high energy $E_{\mathrm{M}}\gg m_{\mathrm{M}}$ in the Higgs boson rest frame. The constituent partons of the vector meson can thus be described by energetic particles moving collinear to the direction of M, and QCD factorization, either in the SCET or NRQCD incarnation, can be employed to compute them.
@@ III.2.1 Higgs decays into a photon plus a vector meson
Table 4 lists the corresponding theoretical predictions and experimental limits for Higgs decays into a photon plus a vector meson, and Fig. 5 displays them in graphical form. Theoretical $\mathcal{B}$ values are in the range of $\mathcal{O}(10^{-5}$$10^{-9})$, with larger rates for decays into light-quark and charm vector mesons compared to bottomonium due to stronger cancellation of direct-indirect contributions for the latter. Experimental upper bounds have been set for all decays at the LHC [ATLAS:2017gko, ATLAS:2018xfc, CMS:2019wch, ATLAS:2023alf, CMS:2024hhg], and HL-LHC could observe $\rm H\to\gamma+\rho$ and $\rm H\to\gamma+\mathrm{J}/\psi$. Most of these decays will be visible at FCC-ee, but heaviest bottomonium radiative decays will require the number of Higgs bosons produced in a machine like FCC-hh.
Table 4: Compilation of exclusive Higgs decays to a photon plus a meson. For each decay, we provide the predicted branching fraction(s) and the theoretical approach used to compute it, as well as the current experimental upper limit and that conservatively estimated for HL-LHC. The last two columns indicate whether the decay can be produced at FCC-ee/FCC-$h$.
@@
$\rm H\to\,\gamma$ ++ $\rm M$ Branching fraction Framework 2024 HL-LHC FCC-$e$ FCC-$h$
[…]
$\left(2.99^{+0.16}_{-0.15}\right)$ $\times~10^{-6}$ NRQCD+LCDA [Bodwin:2017wdu]
$(2.95\pm 0.17)$ $\times~10^{-6}$ SCET+LCDA [Konig:2015qat]
[…]
$\left(9.97^{+4.04}_{-3.03}\right)$ $\times~10^{-9}$ NRQCD (NLL)+LDME [Brambilla:2019fmu]
$\left(5.22^{+2.02}_{-1.70}\right)$ $\times~10^{-9}$ NRQCD+LCDA [Bodwin:2017wdu]
$\left(4.61^{+1.76}_{-1.23}\right)$ $\times~10^{-9}$ SCET+LCDA [Konig:2015qat]
[…]
$\left(2.62^{+1.39}_{-0.91}\right)$ $\times~10^{-9}$ NRQCD (NLL)+LDME [Brambilla:2019fmu]
$\left(1.42^{+0.72}_{-0.57}\right)$ $\times~10^{-9}$ NRQCD+LCDA [Bodwin:2017wdu]
$\left(2.34^{+0.76}_{-1.00}\right)$ $\times~10^{-9}$ SCET+LCDA [Konig:2015qat]
[…]
$\left(1.87^{+1.05}_{-0.69}\right)$ $\times~10^{-9}$ NRQCD (NLL)+LDME [Brambilla:2019fmu]
$\left(9.1^{+4.8}_{-3.8}\right)$ $\times~10^{-10}$ NRQCD+LCDA [Bodwin:2017wdu]
$\left(2.13^{+0.76}_{-1.12}\right)$ $\times~10^{-9}$ SCET+LCDA [Konig:2015qat]
@@ III.2.2 Higgs decays into a Z boson plus a meson
Higgs decays of the form $\rm H\to Z+M$ involve neutral currents and are very similar to the radiative decays just discussed. However, an important difference with respect to the $\gamma\,+\,$meson decays, is that now since the massive final-state gauge boson can be in a longitudinal polarization state, both pseudoscalar and vector mesons can be produced. Also, the interference terms between the direct and indirect amplitudes are smaller. Although these decays are less useful for measuring the Higgs couplings to light quarks, they probe the important effective H-$\gamma$-Z coupling, and provide independent constraints on the meson distribution amplitudes. In addition, they constitute a background for exotic BSM decays of the Higgs boson into, e.g., a Z boson plus an ALP ($\mathrm{H}\to\mathrm{Z}+a$[CMS:2023eos].
@@
Table 5 compiles the theoretical predictions and experimental limits for Higgs decay branching ratios into a Z boson plus a meson, and Fig. 6 presents them in graphical form. All decays have rates in the $10^{-5}$$10^{-6}$ range with small differences across models for the same channel, except for the $\rm H\to Z+\rho$ case where the calculations of [Isidori:2013cla] and [Alte:2016yuw] differ by a factor of six, seemingly because the former had neglected the indirect contributions. Experimental upper bounds have been set for a few channels listed in Table 5, and they are in the $\mathcal{O}(10^{-2}$$10^{-3})$ range [^4] We note that one could a priori also derive limits for $\mathcal{B}(\rm H\to Z+\eta_{\mathrm{c}})$ and $\mathcal{B}(\rm H\to Z+\eta_{\mathrm{b}})$ from a recent 95% CL upper limit on the cross section for the $\rm H\to Z(\ell\ell)a(\gamma\gamma)$ process over $m_{a}\approx 1$–30 GeV performed by the CMS Collaboration [CMS:2023eos]. However, the upper bound cross sections measured at the $m_{a}=m_{\eta_{\mathrm{c}},\eta_{\mathrm{b}}}\approx 2.9,~9.4$ GeV mass points: $\sigma\lesssim 7,$ 2.1 fb, corresponding to $\mathcal{B(\rm H\to Z+a)}\lesssim 0.002,~0.62\cdot 10^{-3}$ after normalizing by the Higgs production cross section and correcting for the dilepton decays of the Z boson) are too weak to place any sensible constrain on these $\mathcal{B}$ (namely, they would exclude the nonphysical region $\mathcal{B}(\rm H\to Z+\eta_{\mathrm{c,b}})\gtrsim 10$, dividing them by the corresponding $\mathcal{B}(\eta_{\mathrm{c,b}}\to\gamma\gamma)$ values).. In general, the experimental $\rm H\to Z+M$ limits are about a factor of ten worse than their $\rm H\to\gamma+M$ counterparts because of the extra events lost from the requirement of Z boson identification via dilepton decays (with $\mathcal{B}\approx 3\%$ for each charged lepton pair). Conservatively, the HL-LHC will be able to set upper bounds about 100 times above their expected SM branching fractions, and it appears that experimental evidence will only be possible at FCC-ee for all of them. Bottomonia-plus-Z and $\eta_{\mathrm{c}}$-plus-Z decays have the largest rates, but no upper-limit has been set to date for them.
@@
Table 5: Compilation of exclusive Higgs decays to a Z boson plus a meson. For each decay, we provide the predicted branching fraction(s) and the theoretical approach used to compute it, as well as the current experimental upper limit and that estimated for HL-LHC. The last two columns indicate whether the decay can be produced at FCC-ee/FCC-$h$.
@@
$\rm H\to\,Z$ ++ $\rm M$ Branching fraction Framework 2024 HL-LHC FCC-$e$ FCC-$h$
$\rm H\,\to\,Z$ + $\rm\pi^{0}$ $(2.3\pm 0.1)$ $\times~10^{-6}$ EFT+NRQM [LHCHiggsCrossSectionWorkingGroup:2016ypw]
$(2.3\pm 0.1)$ $\times~10^{-6}$ EFT+LCDA [Alte:2016yuw]
$\rm\eta$ $(8.3\pm 0.9)$ $\times~10^{-7}$ EFT+LCDA [Alte:2016yuw]
[…]
$(7.19\pm 0.29)$ $\times~10^{-6}$ EFT+LCDA [Alte:2016yuw]
$\rm\omega$ $(1.6\pm 0.1)$ $\times~10^{-6}$ EFT+NRQM [LHCHiggsCrossSectionWorkingGroup:2016ypw]
$(5.6\pm 0.2)$ $\times~10^{-7}$ EFT+LCDA [Alte:2016yuw]
$\rm\eta^{\prime}$ $(1.24\pm 0.13)$ $\times~10^{-6}$ EFT+LCDA [Alte:2016yuw]
[…]
$(2.42\pm 0.10)$ $\times~10^{-6}$ EFT+LCDA [Alte:2016yuw]
$\rm\eta_{c}$ $(1.00\pm 0.01)$ $\times~10^{-5}$ EFT+LCDA [Becirevic:2017chd]
$(1.0\pm 0.1)$ $\times~10^{-5}$ EFT+NRQM [LHCHiggsCrossSectionWorkingGroup:2016ypw]
[…]
$(2.3\pm 0.1)$ $\times~10^{-6}$ EFT+LCDA [Alte:2016yuw]
$3.2$ $\times~10^{-6}$ EFT+NRQM [Gao:2014xlv]
[…]
$\rm\eta_{b}$ $(2.69\pm 0.05)$ $\times~10^{-5}$ EFT+LCDA [Becirevic:2017chd]
$\left(4.739^{+0.276}_{-0.244}\right)$ $\times~10^{-5}$ EFT (NLO)+LCDA [Zhu:2018xxs]
$\rm\Upsilon(1S)$ $1.7$ $\times~10^{-5}$ NRQCD (NLO)+LMDE [Sun:2018xft]
$(1.54\pm 0.06)$ $\times~10^{-5}$ EFT+LCDA [Alte:2016yuw]
$1.7$ $\times~10^{-5}$ EFT+NRQM [Gao:2014xlv]
$\rm\Upsilon(2S)$ $(7.5\pm 0.3)$ $\times~10^{-6}$ EFT+LCDA [Alte:2016yuw]
$8.9$ $\times~10^{-6}$ EFT+NRQM [Gao:2014xlv]
$\rm\Upsilon(3S)$ $(5.63\pm 0.24)$ $\times~10^{-6}$ EFT+LCDA [Alte:2016yuw]
$6.7$ $\times~10^{-6}$ EFT+NRQM [Gao:2014xlv]